OpenStudy (anonymous):

What polynomial has a graph that passes through the given points? (-2, 2) (-1, -1) (1, 5) (3, 67) A. y = -x3 + 4x2 + 2x - 2 B. y = x3 + 4x2 + 2x - 2 C. y = x4 + 4x3 + 2x2 - 2x D. y = x3 - 4x2 - 2x - 2

OpenStudy (aum):

Plug in the x-values and see which one gives the corresponding y-values.

OpenStudy (anonymous):

c?

OpenStudy (aum):

Try the easy point first. That is, (1,5). Put x = 1 in each and see which ones gives you 5

OpenStudy (aum):

After that try (-1, -1).

OpenStudy (anonymous):

ok ok thank you give me one min please

OpenStudy (anonymous):

oh its A? because i got y=5?

OpenStudy (aum):

Try (1,5): A. y = -x3 + 4x2 + 2x - 2 x = 1: -1 + 4 + 2 - 2 = 3 (Not 5. So ruled out) B. y = x3 + 4x2 + 2x - 2 x = 1: 1 + 4 + 2 - 2 = 5 (possibly his one) C. y = x4 + 4x3 + 2x2 - 2x x = 1: 1 + 4 + 2 - 2 = 5 (possibly his one) D. y = x3 - 4x2 - 2x - 2 x = 1: 1 - 4 - 2 - 2 = -7 (Not 5. ruled out)

OpenStudy (anonymous):

b!

OpenStudy (aum):

It is either B or C. Try (-1, -1) B. y = x3 + 4x2 + 2x - 2 x =-1: -1 + 4 - 2 - 2 = -1 (possibly his one) C. y = x4 + 4x3 + 2x2 - 2x x = -1: 1 - 4 + 2 + 2 = 1 (Not -1. ruled out). So it is B.

OpenStudy (anonymous):

What is a quartic function with only the two real zeros given? x = -4 and x = -1 @aum

OpenStudy (anonymous):

x = -4 and x = -1 A. y = x4 + 5x3 + 5x2 + 5x + 4 B. y = x4 - 5x3 - 5x2 - 5x - 4 C. y = -x4 + 5x3 + 5x2 + 5x + 4 D. y = x4 + 5x3 + 5x2 + 5x - 5 \i think d